# Benchmarking Time Series Foundation Models for Short-Term Household Electricity Load Forecasting

### v2

@@ Abstract
1 Introduction
@@
The energy transition, especially the incorporation of renewable energy sources into our energy system, leads to increased electricity load variability, as more households act simultaneously as generators and consumers Tavakoli et al. (2020), electric vehicles introduce additional irregular load into the grid Abid et al. (2024) and the electrical grid is decentralized into micro-grids Shaukat et al. (2023). From a distribution system operator’s (DSO’s) perspective, forecasting the energy consumption of private households poses unique challenges, as their load profiles depend on various (unobserved) factors like household size, installed appliances, or own energy generation (e.g., PV). Therefore, households are often black boxes, leaving reasons for consumption variability unrevealed Paatero and Lund (2006). Additionally, the sheer number of private households leads to massive data management and processing challenges, which require sophisticated machine learning pipelines with continuous training and evaluation of models. Consequently, an efficient and effective electricity demand prediction based on diverse, univariate load time series in the short term, particularly at the low-voltage household level, is vital to ensure a resilient and intelligent electricity distribution.
@@
Various studies have compared univariate low-voltage, household, or residential electricity short-term load forecasting (STLF) approaches Hertel et al. (2023); Mathumitha et al. (2024); Vanting et al. (2021), providing evidence that univariate deep learning approaches based on the Transformer architecture Vaswani et al. (2017) outperform other univariate approaches Hertel et al. (2023); Upadhyay et al. (2023). While Trained-from-Scratch (TFS) Transformers models deliver accurate and fast forecasts, it is necessary to train them from scratch for every specific domain or task (e.g., type of household or geography) and retrain them in regular intervals (e.g., every season).
@@
The recent development of time series foundation models (TSFM), which are (pre-)trained on large and diverse time series datasets, offers the possibility to depart from the traditional method of training one model per task and iteratively retraining it. Out-of-the-box, without further domain adaptation or fine-tuning Ruder (2019), these models can accurately (zero-shot) predict univariate time series from historical data Liang et al. (2024). This advancement could transform the way we forecast household electricity loads by enabling straightforward predictions without continual and task-specific retraining. However, whether massive pre-training of Transformers on very large collections of generic time series (e.g., Finance, Healthcare, Traffic, Energy) Ansari et al. (2024); Rasul et al. (2024) can actually represent household load patterns in real-world scenarios is an empirical question that so far has not been answered, especially considering new evaluation hurdles coming alongside these global models. Accordingly, this study is guided by the central research question:
@@
Our results suggest that TSFM are comparable to TFS Transformers in terms of accuracy. We found that depending on the metric, Time-MoE Shi et al. (2024), Sundial Liu et al. (2025), Chronos Ansari et al. (2024) and TimesFM Das et al. (2024) provide competitive forecasting capabilities. Especially on longer input sizes, TSFM outperformed multiple recent TFS Transformer approaches without being fine-tuned on the task. This finding suggests that with domain adaptation (e.g., creating a foundation model for diverse energy load forecasting tasks) or fine-tuning (e.g., on historical data from the time series under consideration), time series foundation models might become a promising research direction in household electricity STLF.
@@
2 Related Work
Statistical approaches (e.g., SeasonalAverage, ARIMA), as well as more recently machine learning and deep learning approaches based on neural networks, dominate the field of low-voltage level electricity STLF Tzafestas and Tzafestas (2001); Haben et al. (2021); Tarmanini et al. (2023); Vanting et al. (2021). Hopf et al. (2023) conducted a meta-analysis on household electricity STLF and showed that especially hybrid neural networks (NNs) and long short-term memory (LSTM) NNs significantly reduce the forecasting error on the individual (i.e., low-voltage household) level.
Considering deep learning approaches, recent studies show that forecasting methods based on the Transformer Vaswani et al. (2017) architecture tend to outperform other approaches, especially LSTMs, in diverse forecasting domains Ahmed et al. (2023); Lara-Benítez et al. (2021); Li et al. (2023); Nascimento et al. (2023); Sun et al. (2023). Furthermore, various studies compared the performance of different variants of Transformer architectures. Wen et al. (2022) compare different Transformer architectures’ forecasting accuracy using different input lengths and different numbers of layers. While they could not determine a superior architecture, they found that using input sizes exceeding the horizon decreases the forecasting performance of Transformer-based architectures, whereas the performance increases with a rising number of layers Wen et al. (2022). In contrast Gao et al. (2023) found that combining different learning strategies in an adaptive theory-guided framework improves performance compared to the vanilla transformer.
Returning to electricity STLF, many studies focused on electricity STLF at the substation Giacomazzi et al. (2023) or grid level Zhang et al. (2022); Zhao et al. (2021), or investigate multivariate methods Zhang et al. (2021); Cen and Lim (2024), while only four studies focus on the use of TFS Transformers in univariate household electricity STLF (see Table 1).
Table 1: Recent studies on univariate household electricity STLF
@@ 2 Related Work
X X X VanillaTransformer1 Upadhyay et al. (2023)
X X X VanillaTransformer Sievers and Blank (2023)
X X X PatchTST2 Cen and Lim (2024)
X X X X PatchTST Hertel et al. (2023)
X X X X VanillaTransformer Emami et al. (2024)
X X X X X VanillaTransformer3 Saravanan et al. (2024)
X X X X X see Section 4 ours
@@
Upadhyay et al. (2023) proposed a VanillaTransformer with a modified training strategy, which predicts the 25th hour based on the historical 24 hours, while the loss is only computed for the 25th forecast value. They compare their approach with diverse machine learning and deep learning algorithms, including random forests, CNN, and LSTM architectures. Furthermore, they show that the Transformer performs best, directly followed by LSTMs.
@@
The study of Sievers and Blank (2023) compared local, central, and federated learning for CNN, LSTM, and Transformer models. They found that the VanillaTransformer performed best in every training scenario and that local learning, where one model is trained on every dataset, and federated learning, where models are trained locally and then merged into a global model, perform equally well. In contrast, the central learning strategy, where one model is trained using all data on a central server, performed worst Sievers and Blank (2023).
@@
Cen and Lim (2024) developed a modified PatchTST Nie et al. (2023) and compared it to a GRU, LSTM, and multiple other Transformer variants. Their modified PatchTST model outperformed all other approaches, followed by the VanillaTransformer.
@@
While all recent studies incorporate baselines and times series cross-validation, only the study of Hertel et al. (2023) used multiple datasets. They compared diverse Transformer-based approaches with linear regression, an ANN, and an LSTM for forecasting hourly values with a 24-hour, 96-hour, and 720-hour horizon on two datasets. Additionally, they investigated three different training strategies: (1) a local training strategy that trains a separate model for every household, (2) a multivariate strategy that trains one model to predict all households at the same time, and (3) a global strategy that trains one model to forecast multiple households but only one at the same time. In their study, a globally trained PatchTST Nie et al. (2023) model performed best for a 96-hour and 720-hour horizon, and a globally trained VanillaTransformer Vaswani et al. (2017) performed best for a 24-hour horizon.
@@
The comparison of the related work summarized in Table 1 suggests that the VanillaTransformer and PatchTST represent the current SOTA in TFS Transformer architectures. Furthermore, two additional TFS Transformer-based architectures seem promising: The recently published iTransformer model Liu et al. (2024) and the Temporal Fusion Transformer (TFT), which is based on a mixture of LSTM and the attention mechanism Lim et al. (2021) and has shown competitive performance in substation electricity STLF tasks Giacomazzi et al. (2023).
@@
Considering the positive results of recent studies on global forecasting models (Hertel et al. (2023)), the approach to train Transformer-based models on vast amounts of time series data comprising different domains and frequencies as so-called foundation models seems to be a promising avenue for future research on STLF. As general-purpose zero-shot forecasting models, these pre-trained models are able to accurately predict time series without fine-tuning or retraining them on the domain or task-specific datasets Liang et al. (2024).
@@
Two types of foundation models can be distinguished. Large language model-based foundation models, which convert time series into textual representations, sometimes enhanced by other architectures such as Graph Neural Networks, such as PromptCast Xue and Salim (2023), LLMTime Gruver et al. (2023), and FSCA Hu et al. (2025) represent the first type of time series foundation models. However, these models have a high resource utilization and lack scalability and practicability Ansari et al. (2024); Tan et al. (2024). Transformer-based architectures for time series such as TimeGPT-1 Garza and Mergenthaler-Canseco (2023), LagLlama Rasul et al. (2024), Chronos Ansari et al. (2024), and TimesFM Das et al. (2024) comprise the second type of foundation models. Inspired by large language models (LLM), these models are specifically trained on tokenized time series to forecast the most probable token, which encodes an explicit part of a time series.
@@
While these models showed impressive zero-shot capabilities in various domains Aksu et al. (2024), their vast training datasets create a unique evaluation problem. Specifically, many publicly available benchmarking datasets have been used to train these foundation models. Therefore, it is crucial to evaluate them on out-of-sample datasets not included in the foundation models’ training data. In fact, it can quickly happen that the selected evaluation data is already available in the sheer volume of training data. For example, a study evaluated TSFM on the Buildingsbench Dataset Emami et al. (2024), which at first glance appears suitable for an STLF evaluation Saravanan et al. (2024). However, the Buildingsbench Dataset partially bundles other datasets that are already included in many of the TSFM training data, such as the London Smart Meters Dataset (Chronos, Moirai, Time-MoE) or the Portuguese Household Dataset (Chronos, TimesFM, Moirai, Time-MoE). Since information leakage is a potential issue here, the Buildingsbench Dataset is not suitable as an TSFM evaluation dataset for STLF.
@@
Table 2: Time series foundation model training data disclosure
@@
Chronos(-Bolt) X Ansari et al. (2024)
LagLlama X Rasul et al. (2024)
Moirai(-MoE) X Woo et al. (2024); Liu et al. (2024)
Time-MoE X Shi et al. (2024)
Sundial X Liu et al. (2025)
TimeGPT-1 Garza and Mergenthaler-Canseco (2023)
TimesFM (2.0) X Das et al. (2024)
As Table 2 summarizes, most TSFM provide information on their (pre-)training data, which enables an evaluation without test set contamination. In contrast, information about the training data for TimeGPT-1 is not publicly available, disqualifying it from an appropriate evaluation using historical open-source datasets.
It is important to note that the typical evaluation strategy used in TSFM often differs from standard time series cross-validation. Instead of applying a rolling-window or expanding-window validation over the entire time series, TSFMs are usually assessed only on the final observations of the series that corresponds to the forecast horizon Das et al. (2024); Ansari et al. (2024); Liu et al. (2025). In other words, the evaluation does not track how the model performs over time but is instead based on producing a single prediction for the last n time points, with performance aggregated across the collection of time series rather than across multiple time segments.
3 Method and Data
STLF approaches at the industry-, building-, and household-levels have been investigated in diverse studies Haben et al. (2021). While Haben et al. (2021) did not identify an approach that is superior in all situations in their review, they emphasized three important factors for the evaluation of low-voltage electricity STLF approaches: (1) The evaluation must be applied on multiple datasets, (2) it should include appropriate naive and sophisticated baselines, such as SeasonalAverage or deep learning models, and (3) it should apply time series cross-validation. Hence, we developed a data acquisition, model selection, and evaluation strategy that fulfills these requirements.
3.1 Data Acquisition
In Table 3, we show which datasets are used in the pre-training of the TSFM. We first considered a total of nine datasets for benchmarking: The Electricity Dataset Trindade (2015) with hourly electricity consumption from households, shops, and industrial business in Portugal, the Ausgrid Solar Home Dataset Ratnam et al. (2017) with solar energy production and private consumption from clients in Australia, and the London Smart Meters Dataset UKPowerNetworks (2015) from electrical consumption of Households in the United Kingdom are all included in the training set of multiple TSFM.
Table 3: Household Energy Datasets used in foundation model pre-training
@@ 3.1 Data Acquisition
Therefore, they cannot be used for evaluation without the risk of leakage Kapoor and Narayanan (2022) and an overestimation of the performance of the foundation model. The French Household Dataset Hebrail and Berard (2012) contains only a single house, lacking diversity, and the Danish Dataset Energinet (2021) contains energy consumption of whole districts, which is not the focus of our study. Three sourced datasets are not included in any pre-training and can be used for evaluation without any restriction: The Southern Germany Dataset Open Power System Data (2020), the Lower Saxony Dataset Schlemminger et al. (2022) and the REFIT Dataset Murray et al. (2017).
@@
The REFIT Dataset contains household data from the Loughborough area in the United Kingdom Murray et al. (2017) with house characteristics and appliance-by-appliance energy consumption per minute for two years. In this study, we used the hourly aggregated consumption for each of the 20 households.
@@
The Southern Germany Dataset comprises electricity consumption from small businesses and households in the city of Konstanz in Germany Open Power System Data (2020). We filtered the data for households only. Some houses also generate energy from PV panels, and some show consumption from individual devices, such as dishwashers, freezers, heat pumps, and, for one house, an electric vehicle. Keeping the perspective of the DSO, we filtered the data to obtain the grid’s total electricity import from the six households in the dataset.
@@
The Lower Saxony Dataset holds electrical single-family house consumption and partly PV energy generation near the city of Hameln in Germany Schlemminger et al. (2022). We selected the active power for all measured phases. The dataset also includes pumps that are used in the district heating network. These are measured via a separate smart meter and can be treated separately by the DSO and, hence, are not considered in the analysis. Four households with PV systems are dropped due to the lack of a clear separation between grid import, direct usage of solar energy, and grid export, rendering them unsuitable for evaluation of the net energy consumption.
@@
The IDEAL Dataset Pullinger et al. (2021) represents a special case: on the one hand, it contains a large number of households, which significantly increases the informative value of the evaluation. On the other hand, it was used in the Moirai, Sundial and Time-MoE training data. We decided to include the dataset in the evaluation and to exclude the evaluation of the three TSFMs for this dataset. Details can be found in section 3.7. The IDEAL Dataset covers electric, gas, temperature, humidity, and metadata from households in Edinburgh and nearby regions in the UK Pullinger et al. (2021). Some houses contain information about electrical appliances and more detailed information about temperature and gas and heating equipment, as well as weather information. In the study, we use the net load electricity consumption from each of the 254 households.
For all datasets, we used hourly aggregated measurements. The unit of electricity consumption is expressed in kilowatt-hours (kWh). The general information about the evaluation datasets is summarized in Table 4.
Table 4: Information about the datasets

### v3

@@ Abstract
Index Terms:
Household Electricity, Foundation Models, Short-term Load Forecasting, Time Series Transformers
[…]
=-21pt
I Introduction
@@
The energy transition, especially the incorporation of renewable energy sources into our energy system, leads to increased electricity load variability, as more households act simultaneously as generators and consumers [tavakoli_impacts_2020], electric vehicles introduce additional irregular load into the grid [Abid.2024] and the electrical grid is decentralized into micro-grids [Shaukat.2023]. From a distribution system operator’s (DSO’s) perspective, forecasting the energy consumption of private households poses unique challenges, as their load profiles depend on various (unobserved) factors like household size, installed appliances, or own energy generation (e.g., PV). Therefore, households are often black boxes, leaving reasons for consumption variability unrevealed [Paatero.2006]. Additionally, the sheer number of private households leads to massive data management and processing challenges, which require sophisticated machine learning pipelines with continuous training and evaluation of models. Consequently, an efficient and effective electricity demand prediction based on diverse, univariate load time series in the short term, particularly at the low-voltage household level, is vital to ensure a resilient and intelligent electricity distribution.
@@
Various studies have compared univariate low-voltage, household, or residential electricity short-term load forecasting (STLF) approaches [Hertel.2023, Mathumitha.2024, Vanting.2021], providing evidence that univariate deep learning approaches based on the Transformer architecture [Vaswani.2017] outperform other univariate approaches [Hertel.2023, Upadhyay.2023]. While Trained-from-Scratch (TFS) Transformers models deliver accurate and fast forecasts, it is necessary to train them from scratch for every specific domain or task (e.g., type of household or geography) and retrain them in regular intervals (e.g., every season).
@@
The recent development of time series foundation models (TSFM), which are (pre-)trained on large and diverse time series datasets, offers the possibility to depart from the traditional method of training one model per task and iteratively retraining it. Out-of-the-box, without further domain adaptation or fine-tuning [ruder2019neural], these models can accurately (zero-shot) predict univariate time series from historical data [Liang.2024]. This advancement could transform the way we forecast household electricity loads by enabling straightforward predictions without continual and task-specific retraining. However, whether massive pre-training of Transformers on very large collections of generic time series (e.g., Finance, Healthcare, Traffic, Energy) [Ansari.2024, Rasul.2024] can actually represent household load patterns in real-world scenarios is an empirical question that so far has not been answered, especially considering new evaluation hurdles coming alongside these global models. Accordingly, this study is guided by the central research question:
@@
Our results suggest that TSFM are comparable to TFS Transformers in terms of accuracy. We found that depending on the metric, Time-MoE [shi_time-moe_2024], Sundial [liu2025sundial], Chronos [Ansari.2024] and TimesFM [Das2024] provide competitive forecasting capabilities. Especially on longer input sizes, TSFM outperformed multiple recent TFS Transformer approaches without being fine-tuned on the task. This finding suggests that with domain adaptation (e.g., creating a foundation model for diverse energy load forecasting tasks) or fine-tuning (e.g., on historical data from the time series under consideration), time series foundation models might become a promising research direction in household electricity STLF.
@@
II Related Work
Statistical approaches (e.g., SeasonalAverage, ARIMA), as well as more recently machine learning and deep learning approaches based on neural networks, dominate the field of low-voltage level electricity STLF [Tzafestas2001, Haben.2021, Tarmanini.2023, Vanting.2021]. Hopf2023 conducted a meta-analysis on household electricity STLF and showed that especially hybrid neural networks (NNs) and long short-term memory (LSTM) NNs significantly reduce the forecasting error on the individual (i.e., low-voltage household) level.
Considering deep learning approaches, recent studies show that forecasting methods based on the Transformer [Vaswani.2017] architecture tend to outperform other approaches, especially LSTMs, in diverse forecasting domains [Ahmed2023, Lara.2021, Li2023, Nascimento2023, Sun2023]. Furthermore, various studies compared the performance of different variants of Transformer architectures. wen_Transformers_survery_2022 compare different Transformer architectures’ forecasting accuracy using different input lengths and different numbers of layers. While they could not determine a superior architecture, they found that using input sizes exceeding the horizon decreases the forecasting performance of Transformer-based architectures, whereas the performance increases with a rising number of layers [wen_Transformers_survery_2022]. In contrast [GAO2023100142] found that combining different learning strategies in an adaptive theory-guided framework improves performance compared to the vanilla transformer.
Returning to electricity STLF, many studies focused on electricity STLF at the substation [Giacomazzi.2023] or grid level [Zhang.2022, Zhao.2021], or investigate multivariate methods [Zhang.2021, Cen.2024], while only four studies focus on the use of TFS Transformers in univariate household electricity STLF (see Table I).
TABLE I: Recent studies on univariate household electricity STLF
@@ II Related Work
X X X VanillaTransformer1 [Upadhyay.2023]
X X X VanillaTransformer [Sievers.2023]
X X X PatchTST2 [Cen.2024]
X X X X PatchTST [Hertel.2023]
X X X X VanillaTransformer [emami_buildingsbench_2024]
X X X X X VanillaTransformer3 [saravanan_analyzing_2024]
X X X X X see Section IV ours
@@
Upadhyay.2023 proposed a VanillaTransformer with a modified training strategy, which predicts the 25th hour based on the historical 24 hours, while the loss is only computed for the 25th forecast value. They compare their approach with diverse machine learning and deep learning algorithms, including random forests, CNN, and LSTM architectures. Furthermore, they show that the Transformer performs best, directly followed by LSTMs.
@@
The study of Sievers.2023 compared local, central, and federated learning for CNN, LSTM, and Transformer models. They found that the VanillaTransformer performed best in every training scenario and that local learning, where one model is trained on every dataset, and federated learning, where models are trained locally and then merged into a global model, perform equally well. In contrast, the central learning strategy, where one model is trained using all data on a central server, performed worst [Sievers.2023].
@@
Cen.2024 developed a modified PatchTST [nie2023time] and compared it to a GRU, LSTM, and multiple other Transformer variants. Their modified PatchTST model outperformed all other approaches, followed by the VanillaTransformer.
@@
While all recent studies incorporate baselines and times series cross-validation, only the study of Hertel.2023 used multiple datasets. They compared diverse Transformer-based approaches with linear regression, an ANN, and an LSTM for forecasting hourly values with a 24-hour, 96-hour, and 720-hour horizon on two datasets. Additionally, they investigated three different training strategies: (1) a local training strategy that trains a separate model for every household, (2) a multivariate strategy that trains one model to predict all households at the same time, and (3) a global strategy that trains one model to forecast multiple households but only one at the same time. In their study, a globally trained PatchTST [nie2023time] model performed best for a 96-hour and 720-hour horizon, and a globally trained VanillaTransformer [Vaswani.2017] performed best for a 24-hour horizon.
@@
The comparison of the related work summarized in Table I suggests that the VanillaTransformer and PatchTST represent the current SOTA in TFS Transformer architectures. Furthermore, two additional TFS Transformer-based architectures seem promising: The recently published iTransformer model [Liu.2024] and the Temporal Fusion Transformer (TFT), which is based on a mixture of LSTM and the attention mechanism [lim2021temporal] and has shown competitive performance in substation electricity STLF tasks [Giacomazzi.2023].
@@
Considering the positive results of recent studies on global forecasting models ([Hertel.2023]), the approach to train Transformer-based models on vast amounts of time series data comprising different domains and frequencies as so-called foundation models seems to be a promising avenue for future research on STLF. As general-purpose zero-shot forecasting models, these pre-trained models are able to accurately predict time series without fine-tuning or retraining them on the domain or task-specific datasets [Liang.2024].
@@
Two types of foundation models can be distinguished. Large language model-based foundation models, which convert time series into textual representations, sometimes enhanced by other architectures such as Graph Neural Networks, such as PromptCast [Xue.2023], LLMTime [Gruver.2023], and FSCA [hu2025contextalignment] represent the first type of time series foundation models. However, these models have a high resource utilization and lack scalability and practicability [Ansari.2024, Tan2024]. Transformer-based architectures for time series such as TimeGPT-1 [Garza.2023], LagLlama [Rasul.2024], Chronos [Ansari.2024], and TimesFM [Das2024] comprise the second type of foundation models. Inspired by large language models (LLM), these models are specifically trained on tokenized time series to forecast the most probable token, which encodes an explicit part of a time series.
@@
While these models showed impressive zero-shot capabilities in various domains [aksu_gift-eval_2024], their vast training datasets create a unique evaluation problem. Specifically, many publicly available benchmarking datasets have been used to train these foundation models. Therefore, it is crucial to evaluate them on out-of-sample datasets not included in the foundation models’ training data. In fact, it can quickly happen that the selected evaluation data is already available in the sheer volume of training data. For example, a study evaluated TSFM on the Buildingsbench Dataset [emami_buildingsbench_2024], which at first glance appears suitable for an STLF evaluation [saravanan_analyzing_2024]. However, the Buildingsbench Dataset partially bundles other datasets that are already included in many of the TSFM training data, such as the London Smart Meters Dataset (Chronos, Moirai, Time-MoE) or the Portuguese Household Dataset (Chronos, TimesFM, Moirai, Time-MoE). Since information leakage is a potential issue here, the Buildingsbench Dataset is not suitable as an TSFM evaluation dataset for STLF.
As Table II summarizes, most TSFM provide information on their (pre-)training data, which enables an evaluation without test set contamination. In contrast, information about the training data for TimeGPT-1 is not publicly available, disqualifying it from an appropriate evaluation using historical open-source datasets.
@@
TABLE II: Time series foundation model training data disclosure
@@
Chronos(-Bolt) X [Ansari.2024]
LagLlama X [Rasul.2024]
Moirai(-MoE) X [woo_unified_2024, liu_moirai-moe_2024]
Time-MoE X [shi_time-moe_2024]
Sundial X [liu2025sundial]
TimeGPT-1 [Garza.2023]
TimesFM (2.0) X [Das2024]
It is important to note that the typical evaluation strategy used in TSFM often differs from standard time series cross-validation. Instead of applying a rolling-window or expanding-window validation over the entire time series, TSFMs are usually assessed only on the final observations of the series that corresponds to the forecast horizon [Das2024, Ansari.2024, liu2025sundial]. In other words, the evaluation does not track how the model performs over time but is instead based on producing a single prediction for the last n time points, with performance aggregated across the collection of time series rather than across multiple time segments.
III Method and Data
STLF approaches at the industry-, building-, and household-levels have been investigated in diverse studies [Haben.2021]. While Haben.2021 did not identify an approach that is superior in all situations in their review, they emphasized three important factors for the evaluation of low-voltage electricity STLF approaches: (1) The evaluation must be applied on multiple datasets, (2) it should include appropriate naive and sophisticated baselines, such as SeasonalAverage or deep learning models, and (3) it should apply time series cross-validation. Hence, we developed a data acquisition, model selection, and evaluation strategy that fulfills these requirements.
III-A Data Acquisition
TABLE III: Household Energy Datasets used in foundation model pre-training
@@ III-A Data Acquisition
In Table III, we show which datasets are used in the pre-training of the TSFM. We first considered a total of nine datasets for benchmarking: The Electricity Dataset [Trindade2015] with hourly electricity consumption from households, shops, and industrial business in Portugal, the Ausgrid Solar Home Dataset [ratnam_2017] with solar energy production and private consumption from clients in Australia, and the London Smart Meters Dataset [UKPower-2015] from electrical consumption of Households in the United Kingdom are all included in the training set of multiple TSFM.
@@
Therefore, they cannot be used for evaluation without the risk of leakage [kapoor_leakage_2022] and an overestimation of the performance of the foundation model. The French Household Dataset [misc_individual_household_electric_power_consumption_235] contains only a single house, lacking diversity, and the Danish Dataset [energinet-2021] contains energy consumption of whole districts, which is not the focus of our study. Three sourced datasets are not included in any pre-training and can be used for evaluation without any restriction: The Southern Germany Dataset [OPSD_2020], the Lower Saxony Dataset [Schlemminger2022] and the REFIT Dataset [Murray2017].
@@
The REFIT Dataset contains household data from the Loughborough area in the United Kingdom [Murray2017] with house characteristics and appliance-by-appliance energy consumption per minute for two years. In this study, we used the hourly aggregated consumption for each of the 20 households.
@@
The Southern Germany Dataset comprises electricity consumption from small businesses and households in the city of Konstanz in Germany [OPSD_2020]. We filtered the data for households only. Some houses also generate energy from PV panels, and some show consumption from individual devices, such as dishwashers, freezers, heat pumps, and, for one house, an electric vehicle. Keeping the perspective of the DSO, we filtered the data to obtain the grid’s total electricity import from the six households in the dataset. Although this dataset contains only six households, its primary value lies in the extensive duration of data collection rather than the number of unique entities. Since our approach involves the training of TFS Transformers and the application of time-series cross-validation, the temporal depth is the deciding factor for including this dataset. This extended duration allows us to incorporate multiple seasonality patterns into the training while simultaneously yielding a robust test set of 16,959 observations.
@@
The Lower Saxony Dataset holds electrical single-family house consumption and partly PV energy generation near the city of Hameln in Germany [Schlemminger2022]. We selected the active power for all measured phases. The dataset also includes pumps that are used in the district heating network. These are measured via a separate smart meter and can be treated separately by the DSO and, hence, are not considered in the analysis. Furthermore, four households with PV systems were excluded. Their metering configuration did not distinguish clearly between grid import and PV generation (net metering), resulting in ambiguous net-load profiles containing negative values. To ensure a consistent target variable representing household demand, these time series were deemed unsuitable for the evaluation and removed.
@@
The IDEAL Dataset [ideal_2021] represents a special case: on the one hand, it contains a large number of households, which significantly increases the informative value of the evaluation. On the other hand, it was used in the Moirai, Sundial and Time-MoE training data. We decided to include the dataset in the evaluation and to exclude the evaluation of the three TSFMs for this dataset. Details can be found in section III-G. The IDEAL Dataset covers electric, gas, temperature, humidity, and metadata from households in Edinburgh and nearby regions in the UK [ideal_2021]. Some houses contain information about electrical appliances and more detailed information about temperature and gas and heating equipment, as well as weather information. In the study, we use the net load electricity consumption from each of the 254 households.
TABLE IV: Information about the datasets
@@
For all datasets, we used hourly aggregated measurements. The unit of electricity consumption is expressed in kilowatt-hours (kWh). The general information about the evaluation datasets is summarized in Table IV.
III-B Exploratory Data Analysis (EDA)